OpenStudy (anonymous):

-FAN AND MEDAL!!- The science club sold 35 T-shirts last week and 28 T-shirts this week. what is the percentage of decrease in the number of T-shirts sold? 1.20% 2. 25% 3. 80% 4. 75%

OpenStudy (anonymous):

The question is asking you to find the percent change from 35 to 28. This can be represented as a proportion 28/35 28/35 = 80 % But you need to find out the decreased change 35-28 = 7 7/35 = 20 % Does that make sense?

OpenStudy (anonymous):

yes indeed thank you!

OpenStudy (anonymous):

Welcome :)